Two liquids of different densities (ρ1 = 1500 kg/m3, ρ2 = 500 kg/m3) are poured together into a 100-liter tank, filling it. If the resulting density of the mixture is 800 kg/m3, find the respective amounts of liquids used in kg.

m1 =

m2 =


1
Expert's answer
2021-08-16T17:45:18-0400

"\rho_1=1500kg/m^3\\\rho_2=500kg/m^3\\\rho=800kg/m^3\\1liter=0.1m^3"

Weight of mixure ="\rho v"

Weight of mixure ="800\times0.1=80"

"x\rho_1+(0.1-x)\times\rho_2=80"

"x\times1500+(0.1-x)\times500=80"


"1500x+50-500x=80\\1000x=30\\x=0.03"

Weight="80\times9.8=784N"

Mass"(m_1)=0.03\times1000=30kg\\(m_2)=0.07\times1000=70kg"